# Quantum Sensing Using Cold Atoms and Molecules for Nuclear and Particle Physics

### v2

@@ Abstract
Technologies for manipulating single atoms and molecules have advanced drastically in the past decades. Due to their excellent controllability of internal states, atoms and molecules serves as one of the ideal platforms as quantum systems. One major research direction in atomic systems is the precise determination of energy differences between two energy levels in atoms. This field, known as precision measurement, is symbolized by the remarkable fractional uncertainty of or lower achieved in the state-of-the-art atomic clocks. Two-level systems in atoms are sensitive to various external fields and can therefore be function as quantum sensors. The effect of these fields manifests as energy shifts in the two-level system. Traditionally, such shifts are induced by electric or magnetic fields, as recognized even before the advent of the precision spectroscopy with lasers. However, with high-precision measurements, energy shifts can also be caused by hypothetical fields weakly coupled to ordinary matter, or by small effects mediated by massive particles, which are conventionally dealt with in the field of nuclear and particle physics. In most cases, the atomic systems as quantum sensors have not been sensitive enough to detect such effects. Instead, experiments searching for these interactions have placed constraints on coupling constants, except in a few cases where effects are predicted by the Standard Model of particle physics. Nonetheless, measurements and searches for these effects in atomic systems have led to the emergence of a new field of physics. In some cases, they open new parameter spaces to explore in conventionally investigated topics, e.g., dark matter, fifth force, CP-violation, and other physics beyond the Standard Model. In other cases, these measurements provide alternative experimental approaches to established topics, e.g., variations of fundamental constants searched astronomically, parity violation and nuclear structure studied in high-energy scattering experiments, and supersymmetric particles investigated with high-energy colliders. The use of cold atoms and molecules as quantum sensors for phenomena originating from nuclear and particle physics evolved significantly in the past decades. This paper highlights the recent developments in the field.
@@ I Introduction to quantum sensing for nuclear and particle physics
(ii) a measurement of a physical quantity based on the superposition of quantum states.
@@
There are many physical realizations of quantized energy levels defined in (i). Examples are superconducting qubits [2, 3], mechanical harmonic oscillators including micro- or nanoparticles and suspended mirrors [4, 5], spin systems in solid state materials such as nitrogen-vacancy centers [6, 7, 8], cold atoms [9] and molecules [10] that are often trapped, and trapped ions [11, 12]. Among them, this paper focuses on cold atoms, molecules, and ions. Laser cooling and trapping allows them to be kept in the temperature of mK or even lower without any cryogenic systems. Coherence can be maintained for seconds thanks to the excellent isolation from the environment. Rapidly developing trapping and manipulation techniques enable independent and arbitrary control of the quantum state of a single atom, ion, or molecule. This capability makes these systems suitable for quantum sensors.
@@
When an atomic or molecular system is selected as a quantum system, it generally has many energy levels. For the quantum sensing purposes, specific energy levels are typically selected, and control of the states are performed by applying oscillating electromagnetic fields on or near resonance. In such cases, the system can be well approximated as a two-level system consisting of the ground state and the excited state . The relevant quantities are summarized in Fig. 1. The energy difference between the two energy levels are , where is the Planck constant. The excited state has a natural decay linewidth denoted by . The oscillating electromagnetic field has a frequency of detuned by from .
@@
An external field can induce a frequency shift of from in the two-level system. can be converted to the magnitude of the external field, with a known relation between the field strength and . One example is the frequency shift caused by a magnetic field, which is known as the Zeeman effect. When a small external magnetic field is applied, assuming that the ground state is insensitive to the magnetic field, , where is the Bohr magneton, is Landé g-factor for the excited state, and is the magnetic quantum number for the total angular momentum . Frequency shifts can be some other fields. Another well-known effect is the Stark effect, which is induced by an external electric field.
@@
The external field is not necessarily limited to electromagnetic or other well-established fields in conventional physics. Consider a hypothetical bosonic field that induces , whose constituent particle has a mass and thus can be a candidate for dark matter (such fields are discussed in Section VII). If a two-level system is sufficiently sensitive, the field can be detected using this system. Searching for such a hypothetical field effectively serves as a search for new particles. Even if no frequency shift is observed, the measurement sets an upper bound on the interaction strength between the field, meaning that the interaction is weaker than the sensitivity of the two-level system. This falls within the realm of physics beyond Standard Model (BSM), which is one of the major topic in the contemporary particle physics. Another example is the sensitivity of atomic spectroscopy to nuclear charge distribution. The finite size of the nucleus causes a deviation in the energy levels of the electronic states in an atom compared to the case of a point charge. As a result, slight differences in nuclear charge radii between isotopes can be detected as shifts in transition frequencies between isotopes for the same electronic transition, which is called the isotope shift. Here, a structure of fm, which corresponds to an energy scale of MeV, is probed by spectroscopy of electronic excitations at an energy scale of eV.
@@
These two examples illustrate how the quantum sensors, particularly those with precision spectroscopy, are sensitive to the nuclear and particle physics. One approach is to try to detect hypothetical interactions with small couplings to the Standard Model (SM) particles through high-precision measurement, such as state-of-the-art atomic clocks. Examples covered in this paper include searches for time variation of fundamental constants, ultralight dark matter, and new force between an electron and a neutron. Although axions and axion-like particles fall within this type of interaction, they are not discussed in this paper, as their detection schemes do not significantly overlap with the field of cold atoms: instead, vapor cells are commonly used. Some other review papers[13, 14, 15, 16] deal with them in details. Theoretical proposals for these interactions have been tested experimentally, and so far, the experiments give null results, setting constraints on such hypothetical interactions. Another approach involves phenomena conventionally studied through high-energy collisions. In this case, some precision measurement experiments detect high-energy phenomena in atomic systems. The topics covered in this paper are atomic parity violation (APV), electric dipole moment (EDM) searches, and nuclear structure determination via isotope shifts. APV addresses the physics of GeV, corresponding to the mass of the W and Z bosons, whereas EDM searches aim to probe physics of TeV, searching for supersymmetric particles. The effects of massive particles can manifest through loop-diagrams that are higher-order processes than tree-level Feynman diagrams, unless the heavy particle itself mediates the interaction. Note that throughout this paper, the term "higher-order perturbation" specifically refers to such fundamental interactions in particle physics, while the higher-order perturbations of electromagnetic fields are referred to as "higher-order effects."
@@
The effects of nuclear and particle physics phenomena are typically extremely small, compared to the conventional shifts caused by external electromagnetic fields. Distinguishing these tiny shifts from those induced by conventional electromagnetic fields is crucial. In this paper, the term "signal" refers to the quantity aimed to be detected: namely, the effects originating from nuclear and particle physics, e.g., the effect of the field. All other effects that behave similarly to the targeted signal are regarded as "backgrounds". These backgrounds include shifts caused by conventional external fields such as electric and magnetic fields. Even among effects induced by nuclear and particle physics, one can be considered as the signal, while others act as the backgrounds. For example, in searches for BSM physics, other effects from the SM are backgrounds. To discriminate signal from backgrounds properly, multiple methods are employed. In addition, a proper theoretical framework is essential to extract nuclear and particle physics effects from slight energy shifts. The following discussions focus on how to select an appropriate two-level system to extract a desired nuclear and particle physics phenomenon, and how to cancel all backgrounds that can potentially mimic the signal.
@@
Definition (ii) is also broad. In a two-level system, a superposition state is frequently used to detect a phase drift between the two-level system and the driving oscillating electromagnetic field (see Section II). Other methods of utilizing superposition to enhance the sensitivity of quantum sensors are discussed, e.g., in Section IX. Among the three definitions of quantum sensing, Definition (iii) is the most stringent. Entangled states, such as squeezed states and Schrödinger’s cat states, are typically difficult to generate and maintain. As a result, examples of measurements with entangled states are limited. This paper discusses quantum sensing with entanglement when relevant reports are available. Experimental results are covered in Section III, while theoretical proposals are mentioned in Section X.
@@
This paper first briefly summarizes how atomic and molecular systems are prepared as quantum sensors. This section is intended for readers outside of the field of atomic, molecular, and optical physics (AMO) to offer a glimpse of how AMO experiments operate. As a realization of the two-level system, optical atomic clocks are first discussed. This is partly because precision spectroscopy of atomic two-level systems forms the foundation of precision measurements and because optical atomic clocks are the state-of-the-art in this field. Additionally, atomic clocks serve as platforms for exploring nuclear and particle physics through precision measurements. The systematic shifts and uncertainties discussed here provide typical factors to consider in other types of precision spectroscopy. This part includes details of AMO physics. The discussion is followed by searches for variation of fundamental constants. This shows the sensitivity of atomic clocks as quantum sensors, also introducing the importance of two clock architectures: highly charged ion (HCI) clocks and nuclear clocks. These two systems are not only promising candidate for next-generation optical clocks [17, 18], but also serve as quantum sensors for variations of fundamental constants and other nuclear and particle physics phenomena. A related topic is searches for ultralight dark matter, where periodic modulation of fundamental constants are interpreted as potential signals of ultralight dark matter. After this, searches for a new force between an electron and a neutron through isotope shifts are discussed, which also function as precise measurements of the nuclear structure. Then searches for atomic parity violation is mentioned, which involves weak-force interactions between the nucleus and an electron. Finally, the paper discusses searches for EDMs in electrons and nuclei. The section for the variation of fundamental constants, ultralight dark matter, fifth force search with isotope shifts, APV, and EDM searches are structured with separate discussions on the theoretical frameworks and experimental results. These subsections provide overviews of nuclear and particle physics theories and AMO experiments, respectively.
@@
How are these topics related to quantum sensors? As introduced, the basis of quantum sensors discussed in this paper is the two-level system. In an atomic clock, a two-level system is selected as the source of stability. Various systematic shifts, mainly induced by external electromagnetic fields, are suppressed as a clock. This opens the possibility of detecting small signal originating from nuclear and particle physics phenomena. Examples of such signals are the variation of fundamental constants, ultralight dark matter, a new force between an electron and a neutron, and nuclear charge distribution. In the searches for APV, a two-level system with the same parity is chosen as the quantum sensor. The magnitude of parity-violating transition rate between the two-level system quantifies APV as the signal. In EDM searches, a spin 1/2 state serves as the quantum sensor for the EDM, and the deviation of the precession from purely spin system due to the EDM is the signal.
II Typical procedure to utilize atoms and molecules for quantum sensors
@@
In this section, typical experimental procedures for preparing atoms as quantum sensors are briefly summarized. In some cases, for example molecular beam experiments or ion trap experiments, certain cooling and trapping procedures are different from those described below for atoms.
@@
Trapping atoms begins with generating an atomic beam, a process that depends on the vapor pressure of the atoms. For low-vapor-pressure atoms such as Yb or Sr, an atomic oven heated to a few hundred degrees Celsius is used. For alkali atoms, which have relatively high vapor pressure, atomic dispensers are commonly used. The resulting atomic beam has a center velocity of a few hundred m/s, requiring deceleration. A Zeeman slower [19] is sometimes employed for this purpose, where a circularly polarized laser beam slightly off-resonant from a broad atomic transition propagates to the opposite direction of the atomic beam and also a spatially varying magnetic field is applied. This combination slows down atoms within a specific velocity range to m/s.
@@
After being trapped in a MOT, atoms are typically transferred to an optical lattice. An optical lattice is formed by pairs of counter-propagating laser beams that create a standing wave. Due to the ac Stark shift, atoms are attracted to the most (least) intense part of the light field when the trapping laser is red- (blue-) detuned from broad-linewidth transitions [22]. Depending on the number and configuration of the pairs of the counterpropagating laser beams, one-, two-, or three-dimensional lattice can be formed.
For precision measurements, the key role of an optical lattice is to put atoms in a recoil-free environment. Thermal atoms that are not trapped change their motional states upon absorbing or emitting a photon. The energy changes due to this recoil is called recoil energy , where is the mass of the atom, is the speed of light and is the recoil frequency. When the trapping frequency of atoms along the direction of the probe light resonant with the two-level system is , and , the atoms remain in their motional states after absorbing or emitting a photon, making the spectroscopy recoil-free. Intuitively, the momentum of the photon is absorbed by the entire trap, similar to the Mössbauer effect. is called Lamb-Dicke parameter, and when the condition is satisfied, atoms are in Lamb-Dicke regime, where the atoms do not change their motional state when they absorb or emit photons.
In the optical lattice, atoms can be further cooled using various techniques. Demonstrated methods include evaporative cooling [23], sideband-resolved Raman cooling [24, 25], cooling with a narrow-linewidth transition [26], and Sisyphus cooling [27]. The final temperature is very low, sometimes 100 nK or lower. Colder atoms with lower average energy are essential for achieving trapping in a shallow optical lattice typically used in the state-of-the-art optical lattice clocks.
@@
After state preparation is completed, atoms undergo the measurement sequence. While the measurement sequence depends on experiments, there are two basic procedures for a sequence. The sequence is performed on a two-level system as illustrated in Fig. 1. The laser intensity is characterized by the Rabi frequency . A more detailed formalism can be found in standard textbooks on atomic physics [35].
The first sequence is the Rabi sequence, where resonant light drives the two-level system. If the initial state is , the final state after a pulse of duration satisfies . When and , . Such a pulse to drive the system from to is called a pulse. If the pulse duration is half of the pulse, , and this pulse is referred to as a pulse. The relative phase between and in can be controlled by adjusting the phase offset of the driving light field. If is perturbed by external fields, the nominal pulse duration may not result in the desired final state. To mitigate this, composite pulses, which are extensively studied in the field of nuclear magnetic resonance (NMR) [36, 37], can be applied in laser spectroscopy.
@@
To function as a sensor, the final state needs to be detected. The most common detection method involves driving a broad-linewidth transition, e.g., MHz, from using a light field and measuring the response of the atoms. One method is fluorescence imaging, where scattered photons from the atoms are collected by a photodetector or a camera. Alternatively, absorption imaging can be used, in which a camera records the probe light transmitted through the atomic cloud. The presence of atoms causes attenuation of the probe beam due to light scattering in other directions, resulting in a shadow in the image. To obtain the fraction of atoms in over the total atom number, population in also needs to be detected. This is achieved by transferring atoms from back to and imaging them separately, after removing all atoms originally in . This process of transferring atoms from to through another excited state from which atoms decay to is known as repumping.

### v3

# Quantum Sensing Using Atomic Clocks for Nuclear and Particle Physics

@@ Abstract
Technologies for manipulating single atoms have advanced drastically in the past decades. Due to their excellent controllability of internal states, atoms serve as one of the ideal platforms as quantum systems. One major research direction in atomic systems is the precise determination of physical quantities using atoms, which is included in the field of precision measurements. One of such precisely measured physical quantities is energy differences between two energy levels in atoms, which is symbolized by the remarkable fractional uncertainty of or lower achieved in the state-of-the-art atomic clocks. Two-level systems in atoms are sensitive to various external fields and can, therefore, function as quantum sensors. The effect of these fields manifests as energy shifts in the two-level system. Traditionally, such shifts are induced by electric or magnetic fields, as recognized even before the advent of precision spectroscopy with lasers. With high-precision measurements, tiny energy shifts caused by hypothetical fields weakly coupled to ordinary matter or by small effects mediated by massive particles can be potentially detectable, which are conventionally dealt with in the field of nuclear and particle physics. In most cases, the atomic systems as quantum sensors have not been sensitive enough to detect such effects. Instead, experiments searching for these interactions have placed constraints on coupling constants, except in a few cases where effects are predicted by the Standard Model of particle physics. Nonetheless, measurements and searches for these effects in atomic systems have led to the emergence of a new field of physics. In some cases, they open new parameter spaces to explore in conventionally investigated topics, e.g., dark matter, fifth force, and other physics beyond the Standard Model. In other cases, these measurements provide alternative experimental approaches to established topics, e.g., variations of fundamental constants searched for astronomically and nuclear structure studied in high-energy scattering experiments. The use of atomic clocks as quantum sensors for phenomena originating from nuclear and particle physics evolved significantly in the past decades. This paper highlights the recent developments in the field.
@@ I Introduction to quantum sensing for nuclear and particle physics
(ii) a measurement of a physical quantity based on superposition of quantum states.
@@
There are many physical realizations of quantized energy levels defined in (i). Examples are superconducting qubits [2, 3], mechanical harmonic oscillators including micro- or nanoparticles and suspended mirrors [4, 5], spin systems in solid state materials such as nitrogen-vacancy centers [6, 7, 8], cold atoms [9] and molecules [10] that are often trapped, and trapped ions [11, 12]. Among them, this paper focuses on cold atoms and ions. Laser cooling and trapping allows them to be kept at the temperature of mK or lower without any cryogenic systems. Coherence can be maintained for seconds thanks to the excellent isolation from the environment. Rapidly developing trapping and manipulation techniques enable independent and arbitrary control of the quantum state of a single atom or ion. This capability makes these systems suitable for quantum sensors.
@@
When an atomic system is selected as a quantum system, it generally has many energy levels. For quantum sensing purpose, specific energy levels are typically selected, and control of the states is performed by applying oscillating electromagnetic fields on or near resonance. In such cases, the system can be well approximated as a two-level system consisting of the ground state and the excited state . The relevant quantities are summarized in Fig. 1. The energy difference between the two energy levels is , where is the Planck constant. The excited state has a natural decay linewidth denoted by . The oscillating electromagnetic field has a frequency of detuned by from .
@@
An external field can induce a frequency shift of from in the two-level system. can be converted to the magnitude of the external field, with a known relation between the field strength and . One example is the frequency shift caused by a magnetic field, which is known as the Zeeman effect. When a small external magnetic field is applied, assuming that the ground state is insensitive to the magnetic field, , where is the Bohr magneton, is the Landé g-factor for the excited state, and is the magnetic quantum number for the total angular momentum . Frequency shifts can be caused by some other fields. Another well-known effect is the Stark effect, which is induced by an external electric field.
@@
The external field is not necessarily limited to electromagnetic or other well-established fields in conventional physics. Consider a hypothetical bosonic field that induces , whose constituent particle has a mass and thus can be a candidate for dark matter (such fields are discussed in Section VII). If a two-level system is sufficiently sensitive, the field can be detected using this system. Searching for such a hypothetical field effectively serves as a search for new particles. Even if no frequency shift is observed, the measurement sets an upper bound on the interaction strength between the field and the two-level system, meaning that the interaction is weaker than the sensitivity of the two-level system. This falls within the realm of physics beyond Standard Model (BSM), which is one of the major topic in contemporary particle physics. Another example is the sensitivity of atomic spectroscopy to nuclear charge distribution. The finite size of the nucleus causes a deviation in the energy levels of the electronic states in an atom compared to the case of a point charge. As a result, slight differences in nuclear charge radii between isotopes can be detected as shifts in transition frequencies between isotopes for the same electronic transition, which is called the isotope shift. Here, a structure of fm, which corresponds to an energy scale of MeV, is probed by spectroscopy of electronic excitations at an energy scale of eV.
@@
These two examples illustrate how the quantum sensors, particularly those with precision spectroscopy, are sensitive to the nuclear and particle physics. One approach is to try to detect hypothetical interactions with small couplings to the Standard Model (SM) particles through high-precision measurement, such as state-of-the-art atomic clocks. Examples covered in this paper include searches for time variation of fundamental constants, ultralight dark matter, and new force between an electron and a neutron. Although axions and axion-like particles fall within this type of interaction, they are not discussed in this paper. Readers are encouraged to refer to some other review papers [13, 14, 15, 16] that deal with them in detail. Theoretical proposals for these interactions have been tested experimentally, and so far, the experiments give null results, setting constraints on such hypothetical interactions. Another approach involves phenomena conventionally studied through high-energy collisions. In this case, some precision measurement experiments detect high-energy phenomena in atomic systems. The topics covered in this paper are nuclear structure determination via isotope shifts.
@@
The effects of nuclear and particle physics phenomena are typically extremely small, compared to the conventional shifts caused by external electromagnetic fields. Distinguishing these tiny shifts from those induced by conventional electromagnetic fields is crucial. In this paper, the term "signal" refers to the quantity aimed to be detected: namely, the effects originating from nuclear and particle physics, e.g., the effect of the field. All other effects that behave similarly to the targeted signal are regarded as "backgrounds". These backgrounds include shifts caused by conventional external fields such as electric and magnetic fields. Even among effects induced by nuclear and particle physics, one can be considered as the signal, while others act as the backgrounds. For example, in searches for BSM physics, other effects from the SM are backgrounds. To discriminate the signal from backgrounds properly, multiple methods are employed. In addition, a proper theoretical framework is essential to extract nuclear and particle physics effects from slight energy shifts. The following discussions focus on how to select an appropriate two-level system to extract a desired nuclear and particle physics phenomenon, and how to cancel all backgrounds that can potentially mimic the signal.
@@
Definition (ii) is also broad. In a two-level system, a superposition state is frequently used to detect a phase drift between the two-level system and the driving oscillating electromagnetic field (see Section II). Among the three definitions of quantum sensing, Definition (iii) is the most stringent. Entangled states, such as squeezed states and Schrödinger’s cat states, are typically difficult to generate and maintain. As a result, examples of measurements with entangled states are limited. This paper discusses quantum sensing with entanglement when relevant reports are available. Experimental results are covered in Section III.
@@
This paper first briefly summarizes how atomic systems are prepared as quantum sensors. This section is intended for readers outside of the field of atomic, molecular, and optical physics (AMO) to offer a glimpse of how AMO experiments operate. As a realization of the two-level system, optical atomic clocks are first discussed. This is partly because precision spectroscopy of atomic two-level systems forms the foundation of precise determination of the energy difference in the two-level system and because optical atomic clocks are the state-of-the-art in this field. Additionally, atomic clocks serve as a platform for exploring nuclear and particle physics through precision measurements. The systematic shifts and uncertainties discussed here provide typical factors to consider in other types of precision spectroscopy. This part includes details of AMO physics. The discussion is followed by searches for variation of fundamental constants. This shows the sensitivity of atomic clocks as quantum sensors, also introducing the importance of two clock architectures: highly charged ion (HCI) clocks and nuclear clocks. These two systems are not only promising candidate for next-generation optical clocks [17, 18], but also serve as quantum sensors for variations of fundamental constants and other nuclear and particle physics phenomena. A related topic is searches for ultralight dark matter, where periodic modulation of fundamental constants is interpreted as potential signals of ultralight dark matter. After this, searches for a new force between an electron and a neutron through isotope shifts are discussed, which also function as precise measurements of the nuclear structure. The section for the variation of fundamental constants, ultralight dark matter, and fifth force search with isotope shifts are structured with separate discussions on the theoretical frameworks and experimental results. These subsections provide overviews of nuclear and particle physics theories and AMO experiments, respectively.
@@
How are these topics related to quantum sensors? As introduced, the basis of quantum sensors discussed in this paper is the two-level system. In an atomic clock, a two-level system is selected as the source of stability. Various systematic shifts, mainly induced by external electromagnetic fields, are suppressed as a clock. This opens the possibility of detecting small signal originating from nuclear and particle physics phenomena. Examples of such signals are the variation of fundamental constants, ultralight dark matter, a new force between an electron and a neutron, and nuclear charge distribution.
II Typical procedure to utilize atoms for quantum sensors
@@
In this section, typical experimental procedures for preparing atoms as quantum sensors are briefly summarized. In some cases, for example, ion trap experiments, certain cooling and trapping procedures are different from those described below for neutral atoms.
@@
Trapping atoms begins with generating an atomic beam, a process that depends on the vapor pressure of the atoms. For low-vapor-pressure atoms such as Yb or Sr, an atomic oven heated to a few hundred degrees Celsius is used. For alkali atoms, which have relatively high vapor pressure, atomic dispensers are commonly used. The resulting atomic beam has a center velocity of a few hundred m/s, requiring deceleration for trapping. A Zeeman slower [19] is sometimes employed for this purpose, where a circularly polarized laser beam slightly off-resonant from a broad atomic transition propagates to the opposite direction of the atomic beam and a spatially varying magnetic field is also applied. This combination slows down atoms within a specific velocity range to m/s.
@@
After being trapped in a MOT, atoms are typically transferred to an optical lattice. An optical lattice is formed by pairs of counter-propagating laser beams that create standing waves. Due to the ac Stark shift, atoms are attracted to the most (least) intense part of the light field when the trapping laser is red- (blue-) detuned from broad-linewidth transitions [22]. Depending on the number and configuration of the pairs of the counterpropagating laser beams, a one-, two-, or three-dimensional lattice can be formed.
For precision spectroscopy, the key role of an optical lattice is to put atoms in a recoil-free environment. Thermal atoms that are not trapped change their motional states upon absorbing or emitting a photon. The energy changes due to this recoil is called recoil energy , where is the mass of the atom, is the speed of light, and is the recoil frequency. When the trapping frequency of atoms along the direction of the probe light resonant with the two-level system is , and , the atoms remain in their motional states after absorbing or emitting a photon, making the spectroscopy recoil-free. Intuitively, the momentum of the photon is absorbed by the entire trap, similar to the Mössbauer effect. is called the Lamb-Dicke parameter, and when the condition is satisfied, where is the quantum number for for the longitudinal vibrational state, atoms are in Lamb-Dicke regime, where the atoms do not change their motional state when they absorb or emit photons.
In an optical lattice, atoms can be further cooled using various techniques. Demonstrated methods include evaporative cooling [23], sideband-resolved Raman cooling [24, 25], cooling with a narrow-linewidth transition [26], and Sisyphus cooling [27]. The final temperature can be very low, sometimes 100 nK or lower. Colder atoms with lower average energy are essential for achieving trapping in a shallow optical lattice, typically used in the state-of-the-art optical lattice clocks.
@@
After state preparation is completed, atoms undergo a measurement sequence. While the measurement sequence depends on experiments, there are two basic procedures for a sequence. The sequence is performed on a two-level system as illustrated in Fig. 1. The laser intensity is characterized by the Rabi frequency . A more detailed formalism can be found in standard textbooks on atomic physics [35].
The first sequence is the Rabi sequence, where resonant light drives the two-level system. If the initial state is , the final state after a pulse of duration satisfies . When and , . Such a pulse to drive the system from to is called a pulse. If the pulse duration is half of the pulse, , and this pulse is referred to as a pulse. The relative phase between and in can be controlled by adjusting the phase offset of the driving light field. If is perturbed by external fields, the nominal pulse duration may not result in the desired final state. To mitigate this, composite pulses, which are extensively studied in the field of nuclear magnetic resonance [36, 37], can be applied in laser spectroscopy.
@@
To function as a sensor, the final state needs to be detected. The most common detection method involves driving a broad-linewidth transition, e.g., MHz, from using a light field and measuring the response of the atoms. One method is fluorescence imaging, where scattered photons from the atoms are collected by a photodetector or a camera. Alternatively, absorption imaging can be used, in which a camera records the probe light transmitted through the atomic cloud. The presence of atoms causes attenuation of the probe beam due to light scattering in other directions, resulting in a shadow in the image. To obtain the fraction of atoms in over the total atom number, population in also needs to be detected. This is achieved by transferring atoms from back to and imaging them separately, after removing all atoms originally in .